Maria's age is 14 years less than orlandos age. The sum of their ages is 30. Find their ages.
Found 2 solutions by macston, Cromlix:
Answer by macston(5194)

.
M=Maria's age; O=Orlandos's age=M+14yrs
.
M+O=30yrs Substitute for O.
M+M+14yrs=30yrs Subtract 14yrs from each side.
2M=16yrs Divide each side by 2.
M=8yrs ANSWER 1: Maria is 8 years old.
.
O=M+14yrs=8yrs+14rs=22yrs ANSWER 2: Orlando is 22 years old.
.
CHECK:
M+O=30yrs
8yrs+22yrs=30yrs
30yrs=30yrs

Answer by Cromlix(4381)

Hi there,
Maria's age is 14 years less than orlandos age
M = O - 14
Rearrange
O = M + 14
The sum of their ages is 30.
O + M = 30
Replace O with (M + 14) from first equation
M + 14 + M = 30
Collect like terms
2M = 30 - 14
2M = 16
M = 8
Therefore:
Maria is 8 yrs old
Orlando's age is 22 years old
Hope this helps:-)